# Realigning Incentives to Build Better Software: a Holistic Approach to Vendor Accountability

### v1

@@ A.2 European Regulation
The European Union provides an important comparison in regard to software regulation and management. For example, the EU Cybersecurity Strategy [63] includes many acts and directives aimed at governing cybersecurity-related activities and together creates a holistic set of initiatives to incentivize better software development. Some of these regulations are already in effect; some have been put into force, but they provide a grace period for affected stakeholders to comply, and some are still in the draft phase. Below, we present the regulations that directly or indirectly impact (software) liability. A timeline of key regulations and their main components are highlighted in Figure 3.
Figure 3: Timeline of key EU regulations and highlight of their contents (regulations directly affecting software liability and audit in gray).
@@
AI liability. Artificial intelligence has rapidly advanced over the past decade, becoming a key digital technology. In response, the EU introduced the AI Act [44] coming into force in 2024 and drafted the AI Liability Directive (AILD) [16]. The EU recognized that even under the new PLD, victims must prove fault, damage, and causation, especially a difficult task given AI’s complexity and autonomy. Ensuring victims can access compensation is crucial, and AI’s opacity should not hinder justice.
@@
The AILD is still in its draft phase, with the EU finalizing proposed amendments [^19] https://datamatters.sidley.com/2024/03/21/eu-formally-adopts-worlds-first-ai-law/, while the AI Act and the new PLD are already in force. As AI systems involve complex software (and hardware), vendors and stakeholders must take software liability seriously.
Summary. The EU devised an elaborate and strict legal framework, “throwing a hand grenade” [67] on explicit software liability, including AI-enabled software. The situation is quite different in the US.
@@ A.3 US Regulation
As mentioned above, the current U.S. National Cybersecurity Strategy [61] put forward the concept of software liability together with an audit-based liability waiver mechanism. While the significance of this step has not gone unnoticed in both the legal and the tech community, its codification and application are not without challenges.
@@
There seems to be a consensus among scholars that attestation will indeed play an important role [5]. In addition to self-attestation, a cybersecurity “safe harbor”, based on the certification of organizations (with a similar approach already codified in the EU NIS2 Directive [20], has received considerable attention [21]. Proponents argue that a certification safe harbor has the ability (over time) to create incentives to improve collective security practices, including supply chain security, via involving certified upstream technology providers.
Dempsey suggests a hybrid approach, combining a product-based “floor” for liability and a process-based “ceiling” for safe harbor [4]. His proposed mechanism for liability is three-fold. First, a rule-based approach would define the minimum standard of care for software focusing on product features and behaviors to be included or avoided (e.g., any weaknesses tracked in the MITRE Common Weakness Enumeration (CWE) database [^20] https://cwe.mitre.org/). Second, as software can be complex and dynamic, a liability regime and analysis are needed to assess non-trivial design and implementation flaws; essentially, a software audit. Third, to provide a safe harbor to developers who followed secure coding recommendations, the proposal suggests a liability waiver or limitation, which makes the outcome predictable for vendors. Clearly defining the conditions under which a vendor qualifies for safe harbor provides legal certainty, reducing the need for costly and prolonged expert debates over what qualifies as reasonable (non-negligent) conduct [7].
Regulating Open Source Software. The 2021 Executive Order on Improving the Nation’s Cybersecurity [19] catalyzed an initiative in open-source security, leading to the Open Source Software Security Mobilization Plan [68]. This plan outlined ten priority areas, including training, digital signatures, and Software Bill of Materials (SBOM) requirements. A significant legislative step followed with the introduction of the bi-partisan Securing Open Source Software Act [69], which defines the Cybersecurity and Infrastructure Security Agency’s (CISA) role in enhancing open-source security. Arriving in the footsteps of the Log4j debacle [70], the proposal marks a pivotal moment in U.S. policy, recognizing open source as essential to national security and emphasizing the government’s supporting role in its long-term resilience.
@@
As evidenced by continuously emerging critical vulnerabilities in popular open source software, such as CVE-2024-3094212121https://nvd.nist.gov/vuln/detail/CVE-2024-3094 in the Linux compression utility XZ Utils, open source software security, especially in the supply chain context, is crucial. However, we have two additional observations. First, bugs and vulnerabilities in proprietary software could be just as critical. Second, in terms of liability, open source should be excluded (similarly to the EU CRA [23]) not to thwart innovation; however, commercial products built on top of open source libraries should be under scrutiny. Note that the Act itself has not yet been enacted.
@@
Woods proposes lawmakers to strengthen the insurers’ ability to subrogate against software vendors by
@@
Notably, as related to the software lifecycle, first, the EO recognizes the need for security not only in the development of software but also in the delivery of software and application of patches. Second, it also requires federal government vendors to attest to the security of their development practices, providing more (self-reported) transparency. Third, these attestations will be validated by the Cybersecurity and Infrastructure Security Agency (CISA); these validations will be publicly available; therefore, non-government customers will also benefit from the added transparency in their purchasing decisions. What’s more, vendors whose attestations fail validation may be referred to the Department of Justice; this brings similarities with the corresponding EU legislations, such as NIS2. Given its largely bi-partisan nature, the EO has a good chance to stay in effect.

### v2

@@ A.2 European Regulation
The European Union provides an important comparison in regard to software regulation and management. For example, the EU Cybersecurity Strategy [68] includes many acts and directives aimed at governing cybersecurity-related activities and together creates a holistic set of initiatives to incentivize better software development. Some of these regulations are already in effect; some have been put into force, but they provide a grace period for affected stakeholders to comply, and some are still in the draft phase. Below, we present the regulations that directly or indirectly impact (software) liability. A timeline of key regulations and their main components is presented in Figure 3.
Figure 3: Timeline of key EU regulations and highlights of their contents (regulations directly affecting software liability and audit in gray).
@@
AI liability. Artificial intelligence has rapidly advanced over the past decade, becoming a key digital technology. In response, the EU introduced the AI Act [49], coming into force in 2024, and drafted the AI Liability Directive (AILD) [19]. The EU recognized that even under the new PLD, victims must prove fault, damage, and causation, especially a difficult task given AI’s complexity and autonomy. Ensuring victims can access compensation is crucial, and AI’s opacity should not hinder justice.
@@
The AILD has been in a draft phase, with the EU promising the finalization of proposed amendments [^20] https://datamatters.sidley.com/2024/03/21/eu-formally-adopts-worlds-first-ai-law/, while the AI Act and the new PLD are already in force. In a somewhat unexpected turn of events, the EU announced the withdrawal of the AILD in February 2025 together with releasing their 2025 Work Programme. The withdrawal elicited mixed responses with some expressing significant concern while others welcomed legal simplification and targeted deregulation [^21] https://www.euractiv.com/section/tech/news/commission-plans-to-withdraw-ai-liability-directive-draw-mixed-reactions/, enabling AI-based European innovation.
As AI systems involve complex software (and hardware) with AI-specific dependencies, vendors and stakeholders are advised to take software liability seriously.
Summary. The EU devised an elaborate and strict legal framework, “throwing a hand grenade” [72] on explicit software liability, (partially) including AI-enabled software. The situation is quite different in the US.
@@ A.3 US Regulation
As mentioned above, the current U.S. National Cybersecurity Strategy [66] put forward the concept of software liability together with an audit-based liability waiver mechanism. While the significance of this step has not gone unnoticed in both the legal and tech communities, its codification and application are not without challenges.
@@
There seems to be a consensus among scholars that attestation will indeed play an important role [5]. In addition to self-attestation, a cybersecurity “safe harbor”, based on the certification of organizations (with a similar approach already codified in the EU NIS2 Directive [23], has received considerable attention [24]. Proponents argue that a certification safe harbor has the ability (over time) to create incentives to improve collective security practices, including supply chain security, by involving certified upstream technology providers.
Dempsey suggests a hybrid approach, combining a product-based “floor” for liability and a process-based “ceiling” for safe harbor [4]. His proposed mechanism for liability is three-fold. First, a rule-based approach would define the minimum standard of care for software, focusing on product features and behaviors to be included or avoided (e.g., any weaknesses tracked in the MITRE Common Weakness Enumeration (CWE) database [^22] https://cwe.mitre.org/). Second, as software can be complex and dynamic, a liability regime and analysis are needed to assess non-trivial design and implementation flaws; essentially, a software audit. Third, to provide a safe harbor to developers who followed secure coding recommendations, the proposal suggests a liability waiver or limitation, which makes the outcome predictable for vendors. Clearly defining the conditions under which a vendor qualifies for safe harbor provides legal certainty, reducing the need for costly and prolonged expert debates over what qualifies as reasonable (non-negligent) conduct [7].
Regulating Open Source Software. The 2021 Executive Order on Improving the Nation’s Cybersecurity [22] catalyzed an initiative in open-source security, leading to the Open Source Software Security Mobilization Plan [73]. This plan outlined ten priority areas, including training, digital signatures, and Software Bill of Materials (SBOM) requirements. A significant legislative step followed with the introduction of the bipartisan Securing Open Source Software Act [74], which defines the Cybersecurity and Infrastructure Security Agency’s (CISA) role in enhancing open-source security. Arriving in the footsteps of the Log4j debacle [75], the proposal marks a pivotal moment in U.S. policy, recognizing open source as essential to national security and emphasizing the government’s supporting role in its long-term resilience.
@@
As evidenced by continuously emerging critical vulnerabilities in popular open source software, such as CVE-2024-3094232323https://nvd.nist.gov/vuln/detail/CVE-2024-3094 in the Linux compression utility XZ Utils, open source software security, especially in the supply chain context, is crucial. However, we have two additional observations. First, bugs and vulnerabilities in proprietary software could be just as critical. Second, in terms of liability, open source should be excluded (similarly to the EU CRA [26]), not to thwart innovation; however, commercial products built on top of open source libraries should be under scrutiny. Note that the Act itself has not yet been enacted.
@@
Woods proposes that lawmakers strengthen the insurers’ ability to subrogate against software vendors by
@@
Notably, as related to the software lifecycle, first, the EO recognizes the need for security not only in the development of software but also in the delivery of software and the application of patches. Second, it also requires federal government vendors to attest to the security of their development practices, providing more (self-reported) transparency. Third, these attestations will be validated by the Cybersecurity and Infrastructure Security Agency (CISA); these validations will be publicly available; therefore, non-government customers will also benefit from the added transparency in their purchasing decisions. What’s more, vendors whose attestations fail validation may be referred to the Department of Justice; this brings similarities with the corresponding EU legislations, such as NIS2. Given its largely bipartisan nature, the EO has a good chance of staying in effect.